Find the selling price per ounce of mixture of 200 oz. of cologne that costs $5.50 per ounce and 500 oz. of a cologne that costs $2.00 per ounce
Answer by ankor@dixie-net.com(22740)

Find the selling price per ounce of mixture of 200 oz. of cologne that costs $5.50 per ounce and 500 oz. of a cologne that costs $2.00 per ounce
:
A typical mixture problem:
:
We know the total resulting amt will be: 200 + 500 = 700 oz
:
Let x = selling price
:
5.50(200) + 2.00(500) = x(700)
:
1100 + 1000 = 700x
:
700x = 2100
:
x = 2100/700
:
x = $3.00 an oz is the resulting amt
:
:
Check solution
5.5(200) + 2(500) = 3(700)
1100 + 1000 = 2100